# “Sharing, Not Showing Off”: How BeReal Approaches Authentic Self-Presentation on Social Media Through Its Design

### v1

@@ 5.2.3. Combatting toxicity with simplicity
BeReal’s straightforward design, with its focus on authentic photo sharing, significantly reduces toxicity levels compared to platforms like Snapchat, Instagram, and TikTok. The minimal use of comments on BeReal—often the breeding grounds for toxicity on other platforms—was seen by participants as a key factor contributing to a healthier user environment. P20 shared: “People can be pretty rude on other social platforms. But on BeReal, it’s not like Twitter where it’s like you can talk to anybody anytime. You have to find the person, you have to friend them. You have to engage on their posts to be able to contact them. It’s not as loose as maybe Snapchat or Instagram or Twitter.”
@@ 5.3. The Perceived Downsides of BeReal’s Design
Although BeReal has successfully nudged users toward an authenticity norm that mitigates many existing issues associated with mainstream social media platforms, we identified two shortcomings of its design: (1) it is boring and (2), at times, toxic.
@@ 5.3.1. Creating artificial pressure to post
A few participants felt that while the authenticity norm on BeReal carries significance, they felt the application of BeReal’s “rules” (P27) should put user comfort as the priority. As P29 explained: “I think some people are very obsessed with the idea of keeping things real, keeping things authentic…But to me, I think that it’s okay. I think it’s okay to be a little bit less authentic online. Only for reasons though, if you have a reason for it, I think that it’s okay.” They believed it acceptable to deviate from BeReal’s norms to avoid personal discomfort. P09 explained: “If I’m self-conscious about it, I won’t post about it, and I’ll just post about something else.” Similarly, P16 expressed that when they were not in the mood, they would simply point the camera “to my ceiling and then one is at the desk.” The need for positive self-presentation was a particular cause of discomfort. P25 reflected: “I still have the pressure of trying to make myself look cooler on BeReal. So if I’m just sitting alone, just chilling, watching TikToks, I usually don’t post BeReal until I’m out [of] my bed or doing something else at least.”
@@ 5.3.2. Inciting new forms of toxicity
Despite its intentions, BeReal’s simplicity and focus on everyday moments have inadvertently fostered new avenues for toxicity, with users sometimes judging each other’s content as “uninteresting” (P16), “underwhelming” (P19), and “boring” (P26), either expressing these judgments directly to their friends or confidentially with us during the study interviews. While participants were divided over whether the uninteresting nature of the posts stemmed from the posts themselves or the individuals creating them, some participants felt that the lack of engaging content was more directly associated with the users. For instance, P27 remarked, “Some posts show people just lying in bed. I think, ‘That doesn’t seem fun.’ If they do this every day, it appears they have no exciting point in their day.”
Anticipating or encountering such judgments led some participants to feel an amplified fear of negative judgment when posting on BeReal, which usually involved more vulnerability than posts on other social media platforms in general. As P05 noted, “People could make fun of your true self.” This led some participants to “just walk outside and take a picture.” (P29). The participant further explained, “Even if they don’t necessarily go out, they wait and just leave the house a bit, open the door, take a picture, and then just walk back inside.” P20 further pointed out the sometimes paradoxical nature of the authenticity norm on BeReal: “People might judge you negatively because they compare you to the people who aren’t authentic.” Since it is possible that not everyone embraces BeReal’s “authenticity” to the same extent, being authentic could sometimes lead to negative consequences. P25 shared a candid example: “I know for me, we all glamorize those super social party raves type of people who hang out, …Whereas I feel like when I try to be authentic, it’s like I don’t do that and usually it’s just me at home watching my shows. And so, it definitely conflicts with what society expects versus what I want to do with my life.”
@@ 5.3.3. An underwhelming user experience
BeReal’s focus on authenticity ironically results in a dull and monotonous user experience. As P19 noted, the platform’s mandate of authenticity, combined with specific posting times and limited features, feels “overbearing.” Ironically, the platform’s stringent structure and time constraints contribute to repetitiveness and ennui, rendering the overall user experience “underwhelming” (P19) for some participants. Participants also shared that the scarcity of compelling features amplifies the platform’s tedium. BeReal has been likened to a fleeting craze, with users reflecting, “It’s one of those fads, in a sense” (P19), and observing, “I think… BeReal kind of died, and it was just a fad” (P26). It appears that BeReal’s intensely authenticity-driven design cultivates an “underwhelming” user experience, pushing users away from the platform even as it strives to pull them in with its authentic appeal.
@@ 5.4. BeReal’s Potential and Limitations to Support Relationship-Building
Participants also reflected on the extent to which BeReal’s encouragement to share unstaged photos at unexpected times might enable relationship-building (or fail to do so). We examined these reflections from the perspective of Social Penetration Theory and reported on teens’ thoughts about BeReal’s potential for increasing the breadth and depth of self-disclosure.
@@ 5.4.2. Limited potential for deepening relationships
Many of our participants reflected that they want to share different facets of themselves online, suggesting a lack of nuance in the kinds of self-disclosure that BeReal makes possible. They found it limiting that on BeReal they are not able to share “[their] energetic style or [their] talkative style” (P28), that they “like listening to K-pop” (P09), “a screenshot…of a game [they’re] playing or a song [they’re] listening to” (P16), and “thoughts of the day or whatever you’re doing” (P14), for example. They wanted to be able to share “anything from like my camera roll that I took like maybe like ten minutes ago” (P03), or just “a lot of things” (P15) that they would share or see on their Instagram, Finsta in particular, online.
Our participants also shared that what they post on BeReal is “very surface level” (P15) since it is “just a photo” (P05), implying that depth of self-disclosure is also lacking. They expressed that they are not able to share their emotions (P02, P16), personality (P18), values (P27), and political beliefs (P06, P15) as they do on other platforms such as Instagram (P06, P09, P18), Snapchat (P09), or Twitter (P02). They also confirmed that the lack of such varied forms of exchange makes it harder for them to learn more about and feel closer to other users on BeReal than on other platforms where deeper self-disclosures happen. P16, for example, shared: “I think you can’t really communicate your emotions that easily through [BeReal], which to me is a pretty important part of getting to know somebody, what kind of things they’re going through.” Such limitations led participants to suggest BeReal adopt other forms of content-sharing, including live photos or text posts.
@@ 5.4.3. An inability to share with specific others
Though BeReal encourages users to curate their audience, it was difficult for some of our participants to draw that boundary given their offline social dynamics. As Social Penetration Theory explains, people build relationships by disclosing intentionally within the context of a specific relationship. Users’ inability to target who they shared with exacerbated their feeling of intrusion or overwhelmed with making BeReal posts. For example, P17 shared, “I’m in high school, and so a lot of it is just people in my grade that I wouldn’t necessarily trust. …So there have been times where I’ve retaken my BeReal, because I didn’t like the way I looked in a photo. Which I think is not the point or what they originally intended, but I think that in the social situation, like high school is, I feel like that it was bound to happen in a way. And I know people that’ll retake their BeReals 10 times just to get the right angle.” They further emphasized the limits of sharing “authentic” content if they are not close with their BeReal friends, stating, “In my opinion, I think no one is really going to see the authentic you but yourself, or maybe someone you’re really close with, like a partner or a really close friend or really close family member. Because everyone around you isn’t going to be able to see you truly unless they know you really well.”
Similarly, some participants perceived unfiltered photo-sharing to be relatively high-stakes and intimate. For example, P25, who called themself a “homebody”, expressed that they do not want other people to see that they “stay at home all the time.” Such feelings of intrusion were especially evident when participants had friends on BeReal with whom they were not very close. For example, P18 shared, “I do also worry about what I look like, because there’s people that I’m not that close with, that I’m friends with on there, so I care about what I look like, and I cut off my face if I feel like I look bummy.”
@@ 5.4.4. Coerced self-disclosure cannot enable relationship building
Some participants explained that for self-disclosure to be authentic, it must be self-motivated. P29 expressed, “the most important thing is that someone has to be comfortable with [sharing]” and that “if you’re uncomfortable with something, then you’re not being real, per se,” suggesting self-motivation as a prerequisite for authenticity. P19 also shared that BeReal is “technically authentic, but at a price.” They elaborated, “I guess I could call it real, but at the same time, I feel like there’s much more just that isn’t worth it. You get self-conscious, feel the need to do it, you feel like you have to do it.”
Moreover, our participants felt that intentions should underpin the authenticity of self-disclosure. As P19 exemplified, even staged social media photos can be “perfectly authentic” if the user consciously acknowledges their staging. Conversely, an ‘in-the-moment’ BeReal with no staging can be considered inauthentic if the disclosure lacks self-motivation or intentionality. Hence, even in the face of modern social media norms where staging equates to inauthenticity, intentional staging can maintain authenticity in a user-defined sense, even if not communicated to the audience. On a related note, personal significance was also often mentioned as a factor not reflected in BeReal’s platform-defined authenticity norms. P29, for example, reflected: “If you are holding a post and you choose to post later…I think to me is authentic. That shows that this was something important to you…I think it has to mean something to you first, to be authentic.” Conversely, BeReal posts shared without curation were perceived as lacking the depth of self-disclosure necessary for meaningful social connection.
@@ 5.5.1. Bound to fail
Some participants expressed cynicism about the potential for BeReal to achieve its mission of encouraging authentic self-presentation, explaining that the platform cannot escape the fact that on any social media platform, users will seek approval and present overly positive versions of themselves. For example, P17 shared, “Yes, people are showing themselves kind of unedited, but I think also it was kind of bound to fail in some ways. No matter what, people are going to want to impress other people. No matter what, people are going to want to have other people see them look good.” They added “I think any social media platform, no matter what it is, it’s never going to be the authentic …because …if you’re looking for someone else’s approval, that’s not being authentic,” suggesting that authenticity and social media are fundamentally paradoxical. P25 shared a related perspective that online self-disclosure should be approached with caution, stating, “yeah, it is promoting authenticity, but you still have to remind yourself that this is still online, and so your digital footprint is still there, and everything you post will still be there, and people will see that. So don’t get too sucked in. Don’t glamorize it too much because it’s still something that’s on the phone and something. It’s still considered a social media.”
@@ 5.5.2. Overly prescriptive
Some participants outright rejected the idea that they should adhere to BeReal’s nudges, viewing them as either unnecessary pressure or misaligned with their social media usage. P14 epitomized this perspective, stating that BeReal serves primarily as “pure entertainment” and that people use the platform as an “excuse to take a photo and share it.” P25 went a step further, expressing clear discomfort with an overly serious approach to the app’s rules, stating, “If someone…need[s] to do a couple retakes within the two minutes, who gives a crap? Even if it’s not real, but I think everybody’s just taking it too seriously.”
Participants reported that they have plenty of options for flouting the platform’s demands. For example, users can choose to wait before taking photos, and users can technically retake photos infinitely. This flexibility meant that the unconvinced users did not follow BeReal’s norm at all, providing them justification for staging or delaying posts for curated self-presentation. For instance, P09 shared, “Personally, I kind of ignore the fact that they tell me to be more real and I kind of just take a photo whenever I feel like.” P05 echoed this view: “I feel the same as like Instagram because, like, you could honestly post whenever you wanted.”
5.5.3. Nevertheless, a welcomed change
Despite criticisms, the consensus among participants is that BeReal’s initiative to foster authenticity is “a great thing” (P22)—a positive departure from the norm of other social media platforms. For example, P06 stated that they appreciated that BeReal “wants to give authenticity” to its users because “it’s never been done before and so it’s something different to see and …experience.” P29 also shared that they like the “ overall message of what [BeReal is] trying to do” as they feel that BeReal “encourages people to not fall into social media’s expectations for what they should and shouldn’t do, or how their life should and shouldn’t be.” P02 went further, suggesting that “it would be amazing if those apps [i.e., Instagram], or maybe popular people on those apps, would start trends where you were more authentic. So when you got your BeReal, your BeReal notification you would also post something on Instagram to show like vulnerability on all social media apps.” For those who felt dissatisfied with BeReal’s take on “authenticity,” they felt that BeReal is “less inauthentic” (P10n) at the very least, and that it is “the most authentic photo version of yourself” given BeReal photos indeed cannot be edited.
@@ 6. Discussion
Peer relationships play a crucial role in adolescent development, and reciprocal and authentic self-disclosure is necessary for these relationships to develop. Despite the significant role of social media in facilitating such adolescent peer interactions, its capacity to support authentic self-disclosure is often compromised, for example, by context collapse and the pressures of self-presentation. This study delves into BeReal, an app distinguished by its focus on fostering reciprocal, authentic self-presentation.
Our research examines BeReal’s design in supporting authentic self-presentation on social media and assesses its alignment with the developmental and self-presentation needs of teens. Here, we (1) introduce the Social Media Self-Presentation Matrix (SMSM) based on literature to guide the specific design goals in supporting self-presentation online, cognizant of the different goals and processes of self-presentation, (2) discuss the implications of social media design based on the matrix, and (3) propose design recommendations for authentic sharing on social media derived from our study findings.
@@ 6.1.1. Evaluating BeReal’s Design Support for Self-Presentation Using the SMSM
Social media platforms generally promote a Controlled, Self-Enhancing mode of self-presentation, but BeReal has attempted to shift towards a more Automatic, Self-expressive model. This effort has been partially successful due to several design choices, such as but not limited to: randomly timed notifications that discourage staged sharing, daily ephemeral posting that reduces posting pressure, hidden follower counts that support audience curation, and a minimalist design that underscores the platform’s single-minded focus on extemporaneous, unfiltered photo-sharing. These elements collectively fostered user behaviors distinctly different from those on platforms like Instagram and were largely welcomed by users.
@@
However, several factors made this shift challenging. Some users had large friend groups that included people they did not fully trust, making Automatic self-presentation unrealistic. Others felt more self-conscious because the platform nudges them to reveal aspects of their lives they typically choose to conceal, such as social reticence. Further, BeReal restricts users to a single, unfiltered photo at a randomly assigned time with a two-minute window. This demands significant skill for accurate representation relative to self-presentation on other platforms with no such constraints. Furthermore, the general negativity in the perception of social media exemplified by privacy concerns tends to guide self-presentation behavior in a more Controlled direction rather than an Automatic one.

### v2

@@ 5.2.3. Combatting toxicity with simplicity
BeReal’s straightforward design, with its focus on authentic photo sharing, significantly reduces toxicity levels compared to platforms like Snapchat, Instagram, and TikTok. The minimal use of comments on BeReal—often the breeding grounds for toxicity on other platforms—was seen by participants as a key factor contributing to a healthier environment. P20 shared: “People can be pretty rude on other social platforms. But on BeReal, it’s not like Twitter where it’s like you can talk to anybody anytime. You have to find the person, you have to friend them. You have to engage on their posts to be able to contact them. It’s not as loose as maybe Snapchat or Instagram or Twitter.”
@@ 5.3. The Perceived Downsides of BeReal’s Design
Although BeReal has successfully nudged users toward an authenticity norm that mitigates many existing issues associated with mainstream social media platforms, participants raised two shortcomings of BeReal, specifically that it could be both toxic and boring.
@@ 5.3.1. Creating artificial pressure to post
A few participants felt that while the authenticity norm on BeReal carries significance, the application of BeReal’s “rules” (P27) should prioritize user comfort. As P29 explained: “I think some people are very obsessed with the idea of keeping things real, keeping things authentic…But to me, I think that it’s okay. I think it’s okay to be a little bit less authentic online. Only for reasons though, if you have a reason for it, I think that it’s okay.” They believed it acceptable to deviate from BeReal’s norms to avoid personal discomfort. P09 explained: “If I’m self-conscious about it, I won’t post about it, and I’ll just post about something else.” Similarly, P16 expressed that when they were not in the mood, they would simply point the camera “to my ceiling and then one is at the desk.” The need for positive self-presentation was a particular cause of discomfort. P25 reflected: “I still have the pressure of trying to make myself look cooler on BeReal. So if I’m just sitting alone, just chilling, watching TikToks, I usually don’t post BeReal until I’m out [of] my bed or doing something else at least.”
@@ 5.3.2. Inciting new forms of toxicity
Despite its intentions, BeReal’s simplicity and focus on everyday moments inadvertently fostered new avenues for toxicity, with users sometimes judging each other’s content as “uninteresting” (P16), “underwhelming” (P19), and “boring” (P26), either expressing these judgments directly to their friends or confidentially with us during the study interviews. While participants were divided over whether the uninteresting nature of the posts stemmed from the posts themselves or the individuals creating them, some participants perceived that the lack of engaging content was more directly associated with the users. For instance, P27 remarked, “Some posts show people just lying in bed. I think, ‘That doesn’t seem fun.’ If they do this every day, it appears they have no exciting point in their day.”
Anticipating or encountering such judgments led some participants to feel an amplified fear of negative judgment when posting on BeReal, which usually involved more vulnerability than posts on other social media platforms in general. As P05 noted, “People could make fun of your true self.” This led some participants to “just walk outside and take a picture” (P29). The participant further explained, “Even if they don’t necessarily go out, they wait and just leave the house a bit, open the door, take a picture, and then just walk back inside.” P20 further pointed out the sometimes paradoxical nature of the authenticity norm on BeReal: “People might judge you negatively because they compare you to the people who aren’t authentic.” Since it is possible that not everyone embraces BeReal’s “authenticity” to the same extent, being authentic could sometimes lead to negative consequences. P25 shared a candid example: “I know for me, we all glamorize those super social party raves type of people who hang out, …Whereas I feel like when I try to be authentic, it’s like I don’t do that and usually it’s just me at home watching my shows. And so, it definitely conflicts with what society expects versus what I want to do with my life.”
@@ 5.3.3. An underwhelming user experience
BeReal’s focus on authenticity ironically leads to a dull and monotonous user experience. As P19 noted, the platform’s norms around authenticity, combined with specific posting times and limited features, feels overbearing.” Such strict structure and time constraints contribute to repetitiveness and boredom. This renders the overall user experience underwhelming” (P19) for some participants. Some users also viewed BeReal’s popularity as temporary, sharing It’s one of those fads, in a sense” (P19), and observing, I think BeReal kind of died, and it was just a fad” (P26). BeReal’s single-minded focus on authenticity cultivates an “underwhelming” user experience that pushes users away from the platform.
@@ 5.4. BeReal’s Potential and Limitations to Support Relationship-Building
Participants also reflected on the extent to which BeReal’s encouragement to share unstaged photos at unexpected times might enable relationship-building (or fail to do so). We examined these reflections from the perspective of Social Penetration Theory (Altman and Taylor, 1973) and reported on teens’ thoughts about BeReal’s potential for increasing the breadth and depth of self-disclosure.
@@ 5.4.2. Limited potential for deepening relationships
Many of our participants reflected that they want to share different facets of themselves online, suggesting a lack of nuance in the kinds of self-disclosure that BeReal makes possible. They found it limiting that on BeReal they are not able to share “[their] energetic style or [their] talkative style” (P28), that they “like listening to K-pop” (P09), “a screenshot…of a game [they’re] playing or a song [they’re] listening to” (P16), and “thoughts of the day or whatever you’re doing” (P14), for example. They wanted to be able to share “anything from like my camera roll that I took like maybe like ten minutes ago” (P03), or just “a lot of things” (P15) that they would share or see on their social media, Finsta in particular.
Our participants also shared that what they post on BeReal is “very surface level” (P15) since it is “just a photo” (P05), implying that depth of self-disclosure is also lacking. They expressed that they are not able to share their emotions (P02, P16), personality (P18), values (P27), and political beliefs (P06, P15) as they do on other platforms such as Instagram (P06, P09, P18), Snapchat (P09), or Twitter (P02). They also confirmed that the lack of varied forms of exchange makes it harder for them to learn more about and feel closer to other users on BeReal than on other platforms where deeper self-disclosure happens. P16, for example, shared: “I think you can’t really communicate your emotions that easily through [BeReal], which to me is a pretty important part of getting to know somebody, what kind of things they’re going through.” Such limitations led participants to suggest BeReal adopt other forms of content-sharing, including live photos or text posts.
@@ 5.4.3. An inability to share with specific others
Though BeReal encourages users to curate their audience, it was difficult for some of our participants to draw that boundary given their offline social dynamics. As Social Penetration Theory (Altman and Taylor, 1973) explains, people build relationships by disclosing intentionally within the context of a specific relationship. Users’ inability to target who they shared with exacerbated their feeling of intrusion or overwhelmed with making BeReal posts. For example, P17 shared, “I’m in high school, and so a lot of it is just people in my grade that I wouldn’t necessarily trust. …So there have been times where I’ve retaken my BeReal, because I didn’t like the way I looked in a photo. Which I think is not the point or what they originally intended, but I think that in the social situation, like high school is, I feel like that it was bound to happen in a way. And I know people that’ll retake their BeReals 10 times just to get the right angle.” They further emphasized the limits of sharing “authentic” content if they are not close with their BeReal friends, stating, “In my opinion, I think no one is really going to see the authentic you but yourself, or maybe someone you’re really close with, like a partner or a really close friend or really close family member. Because everyone around you isn’t going to be able to see you truly unless they know you really well.”
Similarly, some participants perceived unfiltered photo-sharing to be relatively high-stakes and intimate. For example, P25, who called themself a “homebody,” expressed that they do not want other people to see that they “stay at home all the time.” Such feelings of intrusion were especially evident when participants had friends on BeReal with whom they were not very close. For example, P18 shared, “I do also worry about what I look like, because there’s people that I’m not that close with, that I’m friends with on there, so I care about what I look like, and I cut off my face if I feel like I look bummy.”
@@ 5.4.4. Coerced self-disclosure cannot enable relationship building
Some participants explained that for self-disclosure to be authentic, it must be voluntary and comfortable for the individual. P29 expressed, “the most important thing is that someone has to be comfortable with [sharing]” and that “if you’re uncomfortable with something, then you’re not being real, per se,” suggesting self-motivation as a prerequisite for authenticity. P19 also shared that BeReal is “technically authentic, but at a price.” They elaborated, “I guess I could call it real, but at the same time, I feel like there’s much more just that isn’t worth it. You get self-conscious, feel the need to do it, you feel like you have to do it.”
Our participants emphasized that the authenticity of self-disclosure hinges on the user’s intentions. As P19 exemplified, even staged social media photos can be “perfectly authentic” if the user consciously acknowledges their staging. Conversely, an ‘in-the-moment’ BeReal with no staging can be considered inauthentic if the disclosure lacks self-motivation or intentionality. Contrary to perceptions equating staging with inauthenticity, participants viewed the intentional sharing of personally significant content as authentic. P29, for example, reflected: “If you are holding a post and you choose to post later…I think to me [that] is authentic. That shows that this was something important to you…I think it has to mean something to you first, to be authentic.” Conversely, BeReal posts shared without curation were perceived as lacking the depth of self-disclosure necessary for meaningful social connection.
@@ 5.5.1. Bound to fail
Some participants expressed cynicism about the potential for BeReal to achieve its mission of encouraging authentic self-presentation, explaining that the platform cannot escape the fact that on any social media platform, users will seek approval and present overly positive versions of themselves. For example, P17 shared, “Yes, people are showing themselves kind of unedited, but I think also it was kind of bound to fail in some ways. No matter what, people are going to want to impress other people. No matter what, people are going to want to have other people see them look good.” They added “I think any social media platform, no matter what it is, it’s never going to be the authentic …because …if you’re looking for someone else’s approval, that’s not being authentic,” suggesting that authenticity and social media are fundamentally incompatible. P25 shared a related perspective that online self-disclosure should be approached with caution, stating, “yeah, it is promoting authenticity, but you still have to remind yourself that this is still online, and so your digital footprint is still there, and everything you post will still be there, and people will see that. So don’t get too sucked in. Don’t glamorize it too much because it’s still something that’s on the phone and something. It’s still considered a social media.”
@@ 5.5.2. Overly prescriptive
Some participants outright rejected the idea that they should adhere to BeReal’s prompts, viewing them as either unnecessary pressure or misaligned with their social media usage. P14 epitomized this perspective, stating that BeReal serves primarily as “pure entertainment” and that people use the platform as an “excuse to take a photo and share it.” P25 went a step further, expressing clear discomfort with an overly serious approach to the app’s rules, stating, “If someone…need[s] to do a couple retakes within the two minutes, who gives a crap? Even if it’s not real, but I think everybody’s just taking it too seriously.”
Participants reported various ways to circumvent the platform’s intended use. For example, they can choose to delay taking photos or retake them multiple times. This flexibility allowed unconvinced users to disregard BeReal’s norms entirely, justifying their decision to stage or delay posts for curated self-presentation. P09 illustrated this attitude: “Personally, I kind of ignore the fact that they tell me to be more real and I kind of just take a photo whenever I feel like.” P05 echoed this sentiment, stating: “I feel the same as like Instagram because, like, you could honestly post whenever you wanted.”
5.5.3. Nevertheless, a welcome change
Despite criticisms, the consensus among participants was that BeReal’s initiative to foster authenticity is “a great thing” (P22)—a positive departure from the norm of other broadcast social media platforms. For example, P06 stated that they appreciated that BeReal “wants to give authenticity” to its users because “it’s never been done before and so it’s something different to see and …experience.” P29 also shared that they like the “ overall message of what [BeReal is] trying to do” as they feel that BeReal “encourages people to not fall into social media’s expectations for what they should and shouldn’t do, or how their life should and shouldn’t be.” P02 went further, suggesting that “it would be amazing if those apps [i.e., Instagram], or maybe popular people on those apps, would start trends where you were more authentic. So when you got your BeReal, your BeReal notification you would also post something on Instagram to show like vulnerability on all social media apps.” For those who felt dissatisfied with BeReal’s take on “authenticity,” they felt that BeReal is “less inauthentic” (P10n) at the very least, and that it is “the most authentic photo version of yourself” given BeReal photos cannot be edited.
@@ 6. Discussion
Peer relationships play a crucial role in adolescent development, and reciprocal and authentic self-disclosure is necessary for these relationships to develop. Despite the significant role of social media in facilitating such adolescent peer interactions, its capacity to support authentic self-disclosure is often compromised, for example, by context collapse and the pressures of self-presentation. This study examines BeReal, an app distinguished by its focus on fostering reciprocal, authentic self-presentation.
Our research examines BeReal’s design in supporting authentic self-presentation on social media and assesses its alignment with the developmental and self-presentation needs of teens. Here, we (1) introduce the Social Media Self-Presentation Matrix (SMSM, see Figure 2), a conceptual model of the forms that self-presentation can take, (2) discuss the implications of social media design based on the matrix, and (3) propose design recommendations for authentic sharing on social media derived from our study findings.
@@ 6.1.1. Evaluating BeReal’s Design Support for Self-Presentation Using the SMSM
Social media platforms generally promote a Controlled, Self-Enhancing mode of self-presentation, but BeReal has attempted to shift towards a more Automatic, Self-Expressive model. This effort has been partially successful due to several design choices, such as but not limited to: randomly timed notifications that discourage staged sharing, daily ephemeral posting that reduces posting pressure, hidden follower counts that support audience curation, and a minimalist design that underscores the platform’s single-minded focus on extemporaneous, unfiltered photo-sharing. These elements collectively foster user behaviors differ from those on platforms like Instagram, and these differences were largely welcomed by users.
@@
However, several factors make this shift challenging. Some users have large friend groups that included people they do not fully trust, making Automatic self-presentation unrealistic. Some people feel more self-conscious because the platform encourages them to reveal aspects of their lives they usually prefer to keep private, such as their limited social interactions. Further, BeReal restricts users to a single, unfiltered photo at a randomly assigned time with a two-minute window. This demands significant skill for accurate representation relative to self-presentation on other platforms with no such constraints. Furthermore, the general negativity in the perception of social media exemplified by privacy concerns tends to guide self-presentation behavior in a more Controlled direction rather than an Automatic one.